# Gromov’s Compactness Theorem for the Intrinsic Timed Hausdorff Distance

### v1

@@ 2.6. Review of Timed-Fréchet Maps
Sakovich-Sormani introduced the following notion in [25, Definition 4.21]:
@@
(22) ${\mathcal{N}}=\{x_{1},x_{2},x_{3},...\}\subset X,$
we define a timed-Fréchet map $\kappa_{\tau,X}=\kappa_{\tau,X,{\mathcal{N}}}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$, as follows:
@@
 [^3] In v1 of [25] these were accidentally named timed-Fréchet maps and this will be fixed.
@@
Any timed-Fréchet map, $\kappa_{\tau,X}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty},$ is distance preserving,
@@
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau:X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
@@
(26) $\tau_{\infty}(w_{0},w_{1},w_{2},...)=w_{0},$
endowed with a causal structure defined by
(27) $w\in J^{+}(z)\Longleftrightarrow\tau_{\infty}(w)-\tau_{\infty}(z)=d_{\ell^{\infty}}(w,z),$
then every timed-Fréchet map $\kappa_{\tau,X}:X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$ is time preserving:
@@
(29) $\kappa_{\tau,X}(p)\in J^{+}(\kappa_{\tau,X}(q))\Longleftrightarrow p\in J^{+}(q)$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau:X\to[0,\tau_{max}]$ that encodes causality as in (18).
2.7. Review of the Intrinsic Timed Hausdorff Distance
In [25, Definition 4.22] Sakovich and Sormani introduced the following notion within their definition of the intrinsic timed Hausdorff distance between causally null compactifiable space-times:
@@ 2.7. Review of the Intrinsic Timed Hausdorff Distance
The intrinsic timed Hausdorff distance between two compact timed-metric-spaces, $(X_{i},d_{i},\tau_{i})$, $i=1,2$, is given by
@@
where the infimum is taken over all possible timed-Fréchet maps, $\kappa_{\tau_{1},X_{1}}:X_{1}\to\ell^{\infty}$ and $\kappa_{\tau_{2},X_{2}}:X_{2}\to\ell^{\infty}$ which are found by considering all countable dense collections of points in the $X_{i}$ and all reorderings of these collections of points.
They proved this notion of distance is definite in [25, Theorem 5.14], which we state below.
@@
The intrinsic timed Hausdorff distance between two compact timed-metric-spaces, as in definition 2.9, is definite in the sense that:
(31) $d^{\kappa}_{\tau-H}\Big((X,d_{X},\tau_{X}),(Y,d_{Y},\tau_{Y})\Big)=\inf\,d^{\ell^{\infty}}_{H}(\kappa_{\tau_{X},X}(X),\kappa_{\tau_{Y},Y}(Y))=0$
@@ 3. Proving the Compactness Theorem with Addresses
If $(X_{j},d_{j},\tau_{j})$ are a sequence of compact timed metric spaces that are equibounded,
@@
then a subsequence converges in the intrinsic timed Hausdorff sense to a compact timed metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance and time preserving timed-Fréchet maps,
(36) $\varphi_{j}:X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
@@
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(38) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
@@
(39) $\sup_{\alpha\in{\mathcal{A}}}\|\varphi_{j}({\mathcal{I}}^{j}(\alpha))-\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
@@ 3.1. Selection of Balls in a Compact Metric Space
Recall the definitions of equiboundedness and equicompactness in the statement of Theorem 2.1. The following proposition is from the first part of Gromov’s proof of his Compactness Criterion in [15] that we will apply later to prove our Theorem 3.1.
In the first proposition, the idea is to carefully select a countable dense collection of points covering each metric space and indexing them in a way which keeps track of where each point is located. Part of the statement is taken exactly from Gromov but we add in the precise indexing of the points $x_{a}$ and provide more details in the proof. See Figure 3.1. We will apply this indexing later to define “addresses”.
@@
Suppose $X_{j}$ is a metric space satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$.
Then we can find a countable dense subset
@@
using a uniform index set $A$, not depending on $j$, as described below. For each $i\in{\mathbb{N}}$, we let $A_{i}$ be the finite set,
@@
(44) $p_{i}:A_{i+1}\to A_{i}\qquad p_{i}:(a_{1},\ldots,a_{i+1})\mapsto(a_{1},\ldots,a_{i}).$
Consider the countable set
@@
to be the index set for our countable collection of points with the maps
(46) $I^{j}_{i}:A_{i}\to X_{j}\quad\textrm{ and }\quad I^{j}:A\to X_{j},$
such that
@@
are centers of well chosen balls of radius $\varepsilon_{i}$ in $X_{j}$ for $a\in A_{i}$.
The collection of points
is countably dense in $X_{j}$ and has been carefully indexed to satisfy the following two properties. First, $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
@@
We will complete the proof of the construction and of (50) and (52) by induction on $i\in{\mathbb{N}}$. Note that, once this induction is completed, we will see that the set $I^{j}(A)$ is dense in $X_{j}$ because we can take any $x\in X_{j}$ and any $\delta>0$, and choose $i$ sufficiently large that $\varepsilon_{i}<\delta$, and then apply (50) to see that,
@@
By construction, the set $A$ and its image $I^{j}(A)$ are countable because they are countable unions of finite sets, $A_{i}$, and their images, $I^{j}_{i}(A_{i})$, respectively. Thus we need only complete the inductive proof to prove this proposition.
@@
(61) $p_{1}:A_{2}\to A_{1}\qquad p_{1}:(a_{1},a_{2})\mapsto(a_{1})$
and the map $I^{j}_{2}:A_{2}\to X_{j}$ identifying the center points,
@@ 3.2. The Index Set of Addresses
A natural consequence of our description of Gromov’s Proposition 3.2, is that we have a uniform way of providing “addresses” identifying all points in equicompact collections of metric spaces. See Figure 3.2. We make this precise in the following proposition:
Proposition 3.3.
Given $A$ as in Proposition 3.2, we define an address, $\alpha\in{\mathcal{A}}$, to be a sequence,
@@
(73) $\alpha_{i}=(a_{1},a_{2},...a_{i})\in A_{i}\subset A\qquad\forall i\in{\mathbb{N}}$
so that
@@
(75) $q_{i}:{\mathcal{A}}\to A_{i}\textrm{ s.t. }q_{i}(\alpha)=\alpha_{i}.$
Then for any $X_{j}$ and any maps $I^{j}_{i}\colon A_{i}\to X_{j}$ satisfying the conclusions, (46)- (51), of Proposition 3.2, the map, ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$, defined by
@@
To prove ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$ is well defined, we need only show the sequence of points,
(79) $\{I^{j}_{i}(\alpha_{i}):\,i\in\mathbb{N}\}$
is Cauchy. By $p_{i}(\alpha_{i+1})=\alpha_{i}$, and the properties of $I^{j}_{i}$ proven in Proposition 3.2, we have
@@
So the sequence is Cauchy. Since $X_{j}$ is compact, it is complete and so the limit in (76) exists defining ${\mathcal{I}}^{j}(\alpha)$.
Next we show surjectivity, taking any $x\in X_{j}$. Since $X_{j}$ is covered by balls of radius $\varepsilon_{1}$ we know there exists $(a^{j}_{1}(x))\in A_{1}$ such that
@@ 3.3. Compactness Theorem with Addresses
We can now prove our compactness theorem:
@@
Then we follow the proof of Burago, Burago and Ivanov [5] of Gromov’s Compactness Theorem (to construct a limit space). By the equiboundedness, for any $a,b\in A$, we define
@@
This is symmetric
(91) $d^{j}_{a,b}=d^{j}_{b,a}\qquad\forall a,b\in A,$
and satisfies the triangle inequality
@@
By the equiboundedness of time, for any $a\in A$, we define
@@
By the countability of the index set $A$, we take subsequences of subsequences and diagonalize so that for the final subsequence (which we still denote with $j$), we have for each $a\in A$, convergence of the real numbers:
@@
We take further subsequences and diagonalize further so that for the final subsequence (which we still denote with $j$), we have for each $a,b\in A$,
@@
We now define a quasi metric space,
(98) $(A,d_{\infty})\textrm{ with }d_{\infty}(a,b)=d^{\infty}_{a,b}\in[0,D]$
@@
Take $b]\in X_{\infty}$. Thus, there exist $k\in\mathbb{N}$ such that $b\in A_{k}$. Since for each $j\in\mathbb{N}$ we know that $I_{i}^{j}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, by the pigeon-hole principle, there exists $a\in A_{i}$ such that
@@
Up to here we have followed the proof of Burago, Burago and Ivanov [5] of Gromov’s Compactness Theorem, with corresponding additions to address the fact that we have time functions, to construct a compact timed-metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. Since we will show timed Hausdorff convergence,
@@
(110) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}:X_{j}\to[0,\tau_{max}]\times\prod_{a\in A}[0,D]\subset\ell^{\infty}$
@@
(114) $I^{\infty}_{i}:A_{i}\to X_{\infty}\subset\bar{X}_{\infty}\textrm{ and }I^{\infty}:A\to X_{\infty}\subset\bar{X}_{\infty},$
@@
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (46)-(49), of Proposition 3.2. It satisfies 50 by (106). Moreover, it easily satisfies (51) as follows. Let $a\in A_{i+1}$,
@@
We know that for each $j\in\mathbb{N}$, we know that
@@
Since we have applied Proposition 3.2 for each $X_{j}$, $j\in\mathbb{N}$, and the previous paragraph, for each $j\in\mathbb{N}\sqcup\{\infty\}$ there exist surjective maps, ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}$, $j\in\mathbb{N}\cup\{\infty\}$ defined by
@@
that satisfies the conditions of Proposition 3.3.
@@
Let $\epsilon=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\varepsilon/3$. Then there exists $N_{\epsilon}^{i}\in{\mathbb{N}}$, as in (120), such that for all $j\geq N_{\epsilon}^{i}$ we have (121) for all $j\geq N_{\varepsilon}^{i}$.
@@
(147) $Z=\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\cup\bigcup_{j\in\mathbb{N}}\kappa_{\mathcal{N}_{j}}(X_{j})\subset\prod_{a\in A}[0,D],$
@@
We claim that $Z\subset\ell^{\infty}$ is compact. Let $\{z_{j}\}\subset Z$. If there exists $J\in\mathbb{N}$ such that an infinite number of the $z_{j}$’s is contained in $\kappa_{\mathcal{N}_{J}}(X_{J})$ then by compactness of $\kappa_{\mathcal{N}_{J}}(X_{J})$ there is a further subsequence that converges to a point in $\kappa_{\mathcal{N}_{J}}(X_{J})\subset Z$. Otherwise, assume by passing to a subsequence that $z_{j}\in\kappa_{\mathcal{N}_{j}}(X_{j})$ for all $j\in\mathbb{N}$. By (146), for each $j\in\mathbb{N}$ there exists $w_{j}\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})$ such that

### v2

@@ 2.6. Review of Timed-Fréchet Maps
Sakovich–Sormani introduced the following notion in [26, Definition 4.21]:
@@
(27) ${\mathcal{N}}=\{x_{1},x_{2},x_{3},\ldots\}\subset X,$
we define a timed-Fréchet map [^4] In v1 of [26] these were accidentally named timed-Fréchet maps and this will be fixed. $\kappa_{\tau,X}=\kappa_{\tau,X,{\mathcal{N}}}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$, by
@@
Any timed-Fréchet map, $\kappa_{\tau,X}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty},$ is distance preserving,
@@
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ and any choice of countably dense points in $X$.
@@
(31) $\tau_{\infty}(w_{0},w_{1},w_{2},\ldots)=w_{0},$
and endowed with a causal structure defined by
(32) $\forall z\in[0,\tau_{max}]\times\ell^{\infty}\quad w\in J_{\infty}^{+}(z)\Longleftrightarrow\tau_{\infty}(w)-\tau_{\infty}(z)=d_{\ell^{\infty}}(w,z),$
then every timed-Fréchet map $\kappa_{\tau,X}\colon X\to[0,\tau_{max}]\times\ell^{\infty}\subset\ell^{\infty}$ is time preserving:
@@
(34) $\kappa_{\tau,X}(p)\in J_{\infty}^{+}(\kappa_{\tau,X}(q))\Longleftrightarrow p\in J_{X}^{+}(q)$
for any separable metric space, $(X,d)$, with a Lipschitz one time function, $\tau\colon X\to[0,\tau_{max}]$ that encodes causality as in (23).
2.7. Review of the Intrinsic Timed-Hausdorff Distance
In [26, Definition 4.22] Sakovich and Sormani introduced the following notion within their definition of the intrinsic timed-Hausdorff distance between causally null compactifiable space-times:
@@ 2.7. Review of the Intrinsic Timed-Hausdorff Distance
The intrinsic timed-Hausdorff distance between two compact timed-metric-spaces, $(X_{i},d_{i},\tau_{i})$, $i=1,2$, is given by
@@
where the infimum is taken over all possible timed-Fréchet maps, $\kappa_{\tau_{1},X_{1}}\colon X_{1}\to\ell^{\infty}$ and $\kappa_{\tau_{2},X_{2}}\colon X_{2}\to\ell^{\infty}$ which are found by considering all countable dense collections of points in the $X_{i}$ and all reorderings of these collections of points.
They proved this notion of distance is definite in [26, Theorem 5.14], which we state below.
@@
The intrinsic timed-Hausdorff distance between two compact timed-metric-spaces, as in definition 2.9, is definite in the sense that:
(36) $d^{\kappa}_{\tau-H}\Big((X,d_{X},\tau_{X}),(Y,d_{Y},\tau_{Y})\Big)=0$
@@ 3. Proving the Compactness Theorem with Addresses
If $(X_{j},d_{j},\tau_{j})$ is a sequence of compact timed metric spaces that are equibounded,
@@
then a subsequence converges in the intrinsic timed-Hausdorff sense to a compact timed metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, up to a subsequence, there exist distance and time preserving timed-Fréchet maps,
(41) $\varphi_{j}\colon X_{j}\to[0,\tau_{max}]\times Z\subset\ell^{\infty},$
@@
In addition, there is an uncountable index set $\mathcal{A}$, and surjective index maps,
(43) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
@@
(44) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\varphi_{j}({\mathcal{I}}^{j}(\alpha)),\varphi_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0.$
@@ 3.1. Selection of Balls in a Compact Metric Space
The following proposition is from the first part of Gromov’s proof of his Compactness Criterion in [16]. Since we will apply it later to prove our Theorem 3.1 we provide more details in the proof and add in the precise indexing of the points $x_{a}$. The proposition allows us to carefully select a countable dense collection of points covering each metric space and indexing them in a way which keeps track of where each point is located. See Figure 3.1. We will apply this indexing later to define the index set of “addresses” $\mathcal{A}$.
@@
Suppose $X_{j}$ is a metric space satisfying the equiboundedness and equicompactness conditions as in the statement of Theorem 2.1. Take the sequence $\varepsilon_{i}=2^{-i}$ and let $N_{i}=N(\varepsilon_{i})$ be natural numbers such that $X_{j}$ can be covered by $N_{i}$ balls of radius $\varepsilon_{i}$. Then we can find a countable dense subset
@@
using a uniform index set $A$, not depending on $j$, as described below:
For each $i\in{\mathbb{N}}$, we let $A_{i}$ be the finite set,
@@
(49) $p_{i}\colon A_{i+1}\to A_{i}\qquad p_{i}\colon(a_{1},\ldots,a_{i+1})\mapsto(a_{1},\ldots,a_{i}).$
Then we define the countable index set for our countable collection of points as
@@
We will define maps
(51) $I^{j}_{i}\colon A_{i}\to X_{j}\quad\textrm{ and }\quad I^{j}\colon A\to X_{j},$
where $I^{j}$ is given by
@@
which will be well chosen centers of balls of radius $\varepsilon_{i}$ in $X_{j}$ for $a\in A_{i}$. Specifically, first $I^{j}_{i}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, which means
@@
Hence, the collection of points
is countably dense in $X_{j}$ and has been carefully indexed to satisfy the previous two properties.
@@
We will define $A_{i}$, $p_{i}$ and $I^{j}_{i}$ satisfying (54) and (56) by induction on $i\in{\mathbb{N}}$. Note that, once this induction is completed, we can see that the set $I^{j}(A)$ is dense in $X_{j}$ because we can take any $x\in X_{j}$ and any $\delta>0$, and choose $i$ sufficiently large that $\varepsilon_{i}<\delta$, and then apply (54) to see that,
@@
By construction, the set $A$ and its image $I^{j}(A)$ are countable because they are countable unions of finite sets, $A_{i}$, and their images, $I^{j}_{i}(A_{i})$, respectively. Thus we need only complete the induction to prove this proposition.
@@
(66) $p_{1}\colon A_{2}\to A_{1}\qquad p_{1}\colon(a_{1},a_{2})\mapsto(a_{1})$
and the map $I^{j}_{2}\colon A_{2}\to X_{j}$ identifying the center points,
@@ 3.2. The Index Set of Addresses
A natural consequence of our description of Gromov’s Proposition 3.2, is that we have a uniform way of providing an “addresses” to all points in equicompact collections of metric spaces. More precisely, we have the following definition and proposition.
Definition 3.3.
Given a set $A$ and maps $p_{i}$ as in Proposition 3.2, an address is a sequence
@@
(78) $\alpha_{i}=(a_{1},a_{2},...a_{i})\in A_{i}\subset A\qquad\forall i\in{\mathbb{N}},$
satisfying
@@
We denote by $\mathcal{A}$ the set of all such addresses (see Figure 3.2).
Proposition 3.4.
@@
(80) $q_{i}\colon{\mathcal{A}}\to A_{i}\textrm{ s.t. }q_{i}(\alpha)=\alpha_{i}.$
Then for any $X_{j}$ and any maps $I^{j}_{i}\colon A_{i}\to X_{j}$ satisfying (51)- (55) of Proposition 3.2, the map, ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, defined by
@@
To prove ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$ is well defined, we need only show the sequence of points,
(84) $\{I^{j}_{i}(\alpha_{i})\colon\,i\in\mathbb{N}\}$
is Cauchy. By $p_{i}(\alpha_{i+1})=\alpha_{i}$, and the property (55) for $I^{j}_{i}$ proven in Proposition 3.2, we have
@@
So the sequence is Cauchy. Since $X_{j}$ is compact, it is complete and so the limit in (81) exists, defining ${\mathcal{I}}^{j}(\alpha)$.
Next we show surjectivity. Take any $x\in X_{j}$. We then apply property (54). Since $X_{j}$ is covered by balls of radius $\varepsilon_{1}$ we know there exists $(a^{j}_{1}(x))\in A_{1}$ such that
@@ 3.3. Compactness Theorem with Addresses
We can now prove our compactness theorem.
@@
Then we follow the proof in [5] of Gromov’s Compactness Theorem, to construct a limit space. By the equiboundedness of the distance, for any $a,b\in A$, we define
@@
This definition is symmetric:
(96) $d^{j}_{a,b}=d^{j}_{b,a}\qquad\forall a,b\in A;$
and it satisfies the triangle inequality:
@@
By the equiboundedness of the time functions, for any $a\in A$, we define
@@
By the countability of the index set $A$, we take subsequences of subsequences and diagonalize so that for the final subsequence (which we still index with $j$), we have for each $a\in A$, convergence of the real numbers:
@@
We take further subsequences and diagonalize further, so that for the final subsequence (which we still index with $j$), we have for each $a,b\in A$,
@@
We now define a pseudo-metric space $(A,d_{\infty})$ with
(103) $d_{\infty}(a,b)=d^{\infty}_{a,b}\in[0,D]$
@@
Take $b]\in X_{\infty}$. Thus, there exists $k\in\mathbb{N}$ such that $b\in A_{k}$. Since for each $j\in\mathbb{N}$ we know that $I_{i}^{j}(A_{i})$ is an $\varepsilon_{i}$-net of $X_{j}$, by the pigeon-hole principle, there exists $a\in A_{i}$ such that
@@
Up to here we have followed the proof from [5] of Gromov’s Compactness Theorem, with corresponding additions to address the fact that we have time functions, to construct a compact timed-metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. Since we will show timed-Hausdorff convergence,
@@
(115) $\kappa_{\tau_{j},{\mathcal{N}}_{j}}\colon X_{j}\to[0,\tau_{max}]\times\prod_{a\in A}[0,D]\subset\ell^{\infty}$
@@
(119) $I^{\infty}_{i}\colon A_{i}\to X_{\infty}\subset\bar{X}_{\infty}\quad\textrm{ and }\quad I^{\infty}\colon A\to X_{\infty}\subset\bar{X}_{\infty},$
@@
Note that we have proven that $\bar{X}_{\infty}$ satisfies the conclusions (51)-(57), of Proposition 3.2. It satisfies 54 by (111). Moreover, it easily satisfies (55) as follows. Let $a\in A_{i+1}$,
@@
We know that for each $j\in\mathbb{N}$,
@@
Since we have applied Proposition 3.2 for each $X_{j}$, $j\in\mathbb{N}$, and the previous paragraph, for each $j\in\mathbb{N}\sqcup\{\infty\}$ there exist surjective maps, ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}$, $j\in\mathbb{N}\cup\{\infty\}$ defined by
@@
that satisfy the conditions of Proposition 3.4.
@@
Let $\epsilon=\varepsilon/3$ and $i\in\mathbb{N}$ such that $\varepsilon_{i}<\epsilon$. Then there exists $N_{\epsilon}^{i}\in{\mathbb{N}}$, as in (125), such that for all $j\geq N_{\epsilon}^{i}$ we have (126).
@@
(152) $Z=\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})\cup\bigcup_{j\in\mathbb{N}}\kappa_{\mathcal{N}_{j}}(X_{j})\subset\prod_{a\in A}[0,D]\subset\ell^{\infty},$
@@
We claim that $Z$ is compact. Let $\{z_{j}\}\subset Z$. If there exists $J\in\mathbb{N}$ such that an infinite number of the $z_{j}$’s is contained in $\kappa_{\mathcal{N}_{J}}(X_{J})$ then by compactness of $\kappa_{\mathcal{N}_{J}}(X_{J})$ there is a further subsequence that converges to a point in $\kappa_{\mathcal{N}_{J}}(X_{J})\subset Z$. Otherwise, assume by passing to a subsequence that $z_{j}\in\kappa_{\mathcal{N}_{j}}(X_{j})$ for all $j\in\mathbb{N}$. By (151), for each $j\in\mathbb{N}$ there exists $w_{j}\in\kappa_{\mathcal{N}_{\infty}}(\bar{X}_{\infty})$ such that